OpenStudy (anonymous):

The legs of a right triangle measure 15 and 20. What is the length of the hypotenuse? 25 35 13 10

OpenStudy (igreen):

\(\sf a^2 + b^2 = c^2\) Use the Pythagorean Theorem. Where 'a' and 'b' are the two legs of the right triangle and 'c' is the hypotenuse.

OpenStudy (igreen):

Plug them in and solve for 'c', can you do that? @Jadak

OpenStudy (igreen):

Yep.. or \(\sf c = \sqrt{16^2+20^2}\)

OpenStudy (anonymous):

it would be 15^2 not 16^2 lol

OpenStudy (anonymous):

20 to the power of 2 is 400 +15^2 square root , = 400+225 square root = 625 square root = 25 square root or 25